### v2

@@ 2 Setting and main result
$C^{2}=2\|e^{(0)}\|^{2}+2\|u\|^{2}+8\Lambda^{s}\|u\|_{V_{P_{\rho}}^{s}}+\sigma_{H}^{2}/\Lambda.$ (31)
This is the constant shown in the formulation of Theorem 2.1. Obviously, varying $A$ and $B$ changes the trade-off between initial error, noise variance, and smoothness assumptions in the convergence estimate (28). Note also that $B$ is not part of the algorithm and can be adjusted to any value. This concludes the proof of Theorem 2.1.
@@ 5 Further remarks
$\mathbb{E}^{\prime}(\|\bar{e}^{(m+1)}\|^{2})\leq\alpha_{m}^{2}(\|e^{(m)}\|^{2}+2\Lambda\sigma_{H}^{2}\mu_{m}^{2})+\bar{\alpha}_{m}^{2}(\|u\|^{2}+\mu_{m}^{-1}\|u\|_{V_{P_{\rho}}^{1}}^{2}).$ (32)
@@
Taking into account (29), for $11/2<t<$ 1 we finally arrive at
$\mathbb{E}(\|e^{(m)}\|^{2})\leq\frac{\|e^{(0)}\|^{2}}{(m+1)^{2}}+\frac{2A^{2}\Lambda\sigma_{H}^{2}}{(m+1)^{2t-1}}+\frac{\|u\|^{2}}{m+1}+\frac{A^{-1}\|u\|_{V_{P_{\rho}}^{1}}^{2}}{(m+1)^{1-t}},$ (33)
$m=1,2,\ldots$. This estimate shows more clearly the guaranteed error decay with respect to the initial error $\|e^{(0)}\|^{2}$, the noise variance $\sigma_{H}^{2}$, and the norms $\|u\|^{2}$ and $\|u\|_{V_{P_{\rho}}^{1}}^{2}$ of the solution $u$, respectively, in dependence on $t$. The asymptotically dominant term is of the form
$\mathrm{O}((m+1)^{-\min(2t-1,1-t)})$
and is minimized when $t=2/3$. For this value of $t$ and with $A=(2\Lambda)^{-1}$ we get
@@
There are obvious drawbacks of the whole setting in which Theorem 2.1 is formulated. First, the assumptions are at most qualitative: Since $\mu$, and thus $\rho$, is usually not at our disposal, we cannot verify the assumption $u\in V_{P_{\rho}}^{s}$, nor assess the value of $\sigma_{H}^{2}$. Moreover, although the restriction to learning rates $\mu_{m}$ of the form (6) may not cause problems in view of the results obtained, the choice of optimal values for $t$ and $A$ is by no means obvious. A rule for the adaptive choice of $\mu_{m}$ that does not require knowledge of the values of $s$ and the size of the norms of $u$, but leads to the same quantitative error decay as guaranteed by Theorem 2.1, would be desirable.
@@ 5.2 Difficulties with convergence in $L^{2}_{\rho}(\Omega,Y)$
Our result for the vector-valued case concerned convergence in $V$, which is isometric to the RKHS $H$ generated by $\mathbf{R}$. What we did not succeed in is extending our methods to obtain better asymptotic convergence rates of $f_{u^{(m)}}\to f_{u}$ in the $L^{2}_{\rho}(\Omega,Y)$ norm. Under the assumption (14) about the existence of the minimizer $u$ in (7), error estimates in the $L^{2}_{\rho}(\Omega,Y)$ norm require the study of $\mathbb{E}(\|P_{\rho}^{1/2}e^{(m)}\|^{2})=\mathbb{E}((P_{\rho}e^{(m)},e^{(m)}))$ instead of $\mathbb{E}(\|e^{(m)}\|^{2})$. If, in analogy to (22), one examines the error decomposition
@@
$+$ $\alpha_{m}^{2}(\|P_{\rho}^{1/2}e^{(m)}\|^{2}-2\mu_{m}(P_{\rho}e^{(m)},P_{\omega_{m}}e^{(m)}))+\mu_{m}^{2}\|P_{\rho}^{1/2}P_{\omega_{m}}e^{(m)}\|^{2}).$
@@
$-2\alpha_{m}^{2}\mu_{m}\|P_{\rho}e^{(m)}\|^{2}$
on the right-hand side which must compensate for positive contributions from terms such as
@@
Since in general $P_{\rho}$ does not commute with the operators $P_{\omega}$, this strategy does not work without additional assumptions.
@@ 5.3 A divergence result
$g:=\mathbb{E}(R_{\omega}y)=\mathbb{E}(R_{\omega}f_{\mu}(\omega))\not\in\mathrm{ran}(P_{\rho}).$ (35)
then the sequence $u^{(m)}$ obtained from the online algorithm (4)-(6) diverges in expectation to $\infty$ in $V$ for any choice of the parameters $11/2<t<$ 1 and $0<A\leq(2\Lambda)^{-1}$. This negative result is equivalent to proving
@@
[…]
To establish (36), we expand $g$ and $u^{(0)}$ with respect to the CONS $\Psi$, and derive an inhomogeneous linear recursion for the coefficients of the expected error trajectory $U^{(m)}:=\mathbb{E}(u^{(m)})$. To do this, set
@@
$U^{(m+1)}=\alpha_{m}(U^{(m)}+\mu_{m}(g-P_{\rho}U^{(m)}))=\alpha_{m}(I-\mu_{m}P_{\rho})U^{(m)}+\alpha_{m}\mu_{m}g,\quad m\geq 0$
@@
m=1,2,…. By linearity of the expectation operator and the fact that $\Psi$ consists of the eigenvectors of $P_{\rho}$, we get
@@
$m=1,2,\ldots$, separately for each $i\in\mathbb{N}$. Under the restrictions on the parameters $\mu_{m}$ from (6), we have $1/2\leq 1-\mu_{l}\lambda_{i}\leq 1$. This shows that we have $0<\epsilon_{m}(\lambda_{i})\leq 1$ for the factor in front of $c_{i}$, which implies that the contribution of the initial guess $u^{(0)}$ can be neglected for $m\to\infty$.
@@
$\geq$ $\frac{c_{0}^{2}}{2}\left(\sum_{\lambda_{i}\geq c_{0}(m+1)^{t-1}}\lambda_{i}^{-2}g_{i}^{2}\right)-\frac{\|u^{(0)}\|^{2}}{(m+1)^{2}},$
@@
It remains to show (38). Set
$\Pi_{k}^{m-1}(\lambda):=\prod_{l=k}^{m-1}(1-\mu_{k}\lambda),\quad k=0,\ldots,m-1,\qquad\Pi_{m}^{m-1}(\lambda)=1.$
@@
$d_{m}(\lambda)\geq\frac{1}{2}\sum_{k\in I_{m}(\lambda)}k\mu_{k-1},\qquad I_{m}(\lambda)=\{k\leq m:\;C\lambda((m+1)^{1-t}-(k+1)^{1-t})\leq 1\}.$ (39)
@@
Substituting this, $k\geq k_{0}+1\geq c_{2}(m+1)$, and $\mu_{k-1}=Ak^{-t}$ into (39) we get
@@ 5.4 A special case
$y_{m}=(u,\psi_{i_{m}})+\epsilon_{m},\qquad m=0,1,\ldots$ (40)
@@
where the coefficients $\alpha_{m}$ and $\mu_{m}$ are given by (6) with $\Lambda=1$. This is a special instance of (4) if we set $\Omega=\mathbb{N}$, $Y=\mathbb{R}$ and define $R_{i}:\,\mathbb{R}\to V$ and $R^{\ast}_{i}:\,V\to\mathbb{R}$ by $R_{i}y=y\psi_{i}$ and $R_{i}^{\ast}v=(v,\psi_{i})$, respectively. To simplify things further, let $i_{m}$ be i.i.d. samples from $\mathbb{N}$ with respect to a discrete probability measure $\rho$ on $\mathbb{N}$, and let $\epsilon_{m}$ be i.i.d. random noise with zero mean and finite variance $\sigma^{2}<\infty$ that is independent of $i_{m}$. The corresponding operator $P_{\rho}$ is given by
@@
its eigenvalues $\lambda_{i}=\rho_{i}$ are given by $\rho$, and it is trace class (w.l.o.g., we assume $\rho_{1}\geq\rho_{2}\geq\ldots$). The spaces $V_{P_{\rho}}^{s}$, $-\infty<s<\infty$, can now be identified as sets of formal orthogonal series
@@
and, as expected, has $u$ as its unique solution. This example also shows that sometimes it is more important to consider convergence in $V$ than convergence in the sense of $L_{\rho}^{2}(\Omega,Y)$.
The simplicity of this example allows a rather comprehensive convergence theory with respect to the scale of $V_{P_{\rho}}^{s}$ spaces. We state the following results without detailed proof.
@@
Let $\,-1\leq\bar{s}\leq 0\leq s$, and $\bar{s}<s\leq\bar{s}+2$. Then, for the sampling process described above, the online algorithm (41) converges for $u\in V_{P_{\rho}}^{s}$ in the $V_{P_{\rho}}^{\bar{s}}$ norm with the bound
$\mathbb{E}(\|e^{(m)}\|^{2}_{V_{P_{\rho}}^{\bar{s}}})\leq C(m+1)^{-\min(\frac{s-\bar{s}}{s+2},\frac{2}{\bar{s}+4})}(A^{\bar{s}-s}\|u\|^{2}_{V_{P_{\rho}}^{s}}+A^{2+\bar{s}}\sigma^{2}),\quad m=1,2,\ldots,$ (42)
if the parameters $t$ and $A$ in (6) satisfy
@@
Let $-1\leq\bar{s}\leq 0\leq s$, $\bar{s}<s$ and $\sigma>0$. For the online algorithm (41) we have
$\sup_{\rho}\sup_{u:\,\|u\|_{V_{P_{\rho}}^{s}}=1}(m+1)^{\min((s-\bar{s})/(2+s),2/(\bar{s}+4))}\mathbb{E}(\|e^{(m)}\|^{2}_{V_{P_{\rho}}^{\bar{s}}})\geq c>0,$ (43)
@@
$\Pi_{k}^{m-1}=(1-\frac{a}{m^{t}})\cdot\ldots(1-\frac{a}{(k+1)^{t}}),\quad 0\leq k\leq m-1,\quad\Pi_{m}^{m-1}=1,$ (46)
@@
A matching lower bound for $\varepsilon_{m,i}$ can be obtained by using a slightly different value of $a$ in the definition of the products $\Pi_{k}^{m-1}$. The remainder of the argument for Theorem 5.1 first requires the substitution of tight upper bounds for $\Pi_{k}^{m-1}$ and $S_{k}$ in dependence on $t$ and $a$ into the bounds for $\varepsilon_{m,i}$. Next, after substituting the estimates for $\varepsilon_{m,i}$ into (44), the resulting series has to be estimated separately for the index sets $I_{1}:=\{i:\,A\rho_{i}\leq(m+1)^{t-1}\}$ and $I_{2}:=\mathbb{N}\backslash I_{1}$ followed by choosing the indicated optimal value of $t=t_{s,\bar{s}}$. This leads to the bound (42) in Theorem 5.1. For the proof of Theorem 5.2, lower bounds for $\Pi_{k}^{m-1}$, $S_{k}$ and consequently for $\varepsilon_{m}$ are needed, combined with choosing suitable discrete probability distributions $\rho$. Regarding lower bounds for $\Pi_{k}^{m-1}$, see the considerations in Subsection 5.3.

### v5

@@ 2 Setting and main result
$C^{2}=2\|e^{(0)}\|^{2}+2\|u\|^{2}+8\Lambda^{s}\|u\|_{V_{P_{\rho}}^{s}}+\sigma_{H}^{2}/\Lambda.$
This is the constant shown in the formulation of Theorem 2.1. Obviously, varying $A$ and $B$ changes the trade-off between initial error, noise variance, and smoothness assumptions in the convergence estimate (23). Note also that $B$ is not part of the algorithm and can be adjusted to any value. Finally, the analysis of (25) shows that the bound (13) holds with some constant $C$ and the exponent $s/(s+2)$ replaced by $\min(2t-1,(1-t)s)$ for arbitrary $1/2<t<1$ and $0<A\leq(2\Lambda)^{-1}$ in (5), see Subsection 5.1 for details in the case $s=1$. This concludes the proof of Theorem 2.1.
@@ 5 Further remarks
$\mathbb{E}^{\prime}(\|\bar{e}^{(m+1)}\|^{2})\leq\alpha_{m}^{2}(\|e^{(m)}\|^{2}+2\Lambda\sigma_{H}^{2}\mu_{m}^{2})+\bar{\alpha}_{m}^{2}(\|u\|^{2}+\mu_{m}^{-1}\|u\|_{V_{P_{\rho}}^{1}}^{2}).$
@@
Taking (24) into account, we finally arrive for $1/2<t<1$ at
$\mathbb{E}(\|e^{(m)}\|^{2})\leq\frac{\|e^{(0)}\|^{2}}{(m+1)^{2}}+\frac{2A^{2}\Lambda\sigma_{H}^{2}}{(m+1)^{2t-1}}+\frac{\|u\|^{2}}{m+1}+\frac{A^{-1}\|u\|_{V_{P_{\rho}}^{1}}^{2}}{(m+1)^{1-t}},$
$m=1,2,\ldots$. This estimate shows more clearly the guaranteed error decay with respect to the initial error $\|e^{(0)}\|^{2}$, the noise variance $\sigma_{H}^{2}$, and the norms $\|u\|^{2}$ and $\|u\|_{V_{P_{\rho}}^{1}}^{2}$ of the solution $u$ in dependence on $t$. The asymptotically dominant term is of the form $\mathrm{O}((m+1)^{-\min(2t-1,1-t)})$ and is minimized when $t=2/3$. For this value of $t$ and with $A=(2\Lambda)^{-1}$ we get
@@
There are obvious drawbacks of the whole setting in which Theorem 2.1 is formulated. First, the assumptions are at most qualitative: Since $\mu$, and thus $\rho$, is usually not at our disposal, we cannot verify the assumption $u\in V_{P_{\rho}}^{s}$, nor assess the value of $\sigma_{H}^{2}$. Moreover, although the restriction to learning rates $\mu_{m}$ of the form (5) may not cause problems in view of the results obtained, the choice of optimal values for $t$ and $A$ is by no means obvious. It would be desirable to have a rule for the adaptive choice of $\mu_{m}$ that does not require knowledge of the values of $s$ and the size of the norms of $u$, but leads to the same quantitative error decay as guaranteed by Theorem 2.1.
@@ 5.2 Difficulties with convergence in $L^{2}_{\rho}(\Omega,Y)$
Our result for the vector-valued case concerned convergence in $V$, which is identical with the RKHS $H$ generated by $\mathbf{R}$. What we did not succeed in is extending our methods to obtain better asymptotic convergence rates of $f_{u^{(m)}}\to f_{u}$ in the $L^{2}_{\rho}(\Omega,Y)$ norm. Under the assumption (10) about the existence of the minimizer $u$ in (6), error estimates in the $L^{2}_{\rho}(\Omega,Y)$ norm require the study of $\mathbb{E}(\|P_{\rho}^{1/2}e^{(m)}\|^{2})=\mathbb{E}((P_{\rho}e^{(m)},e^{(m)}))$ instead of $\mathbb{E}(\|e^{(m)}\|^{2})$. If, in analogy to (17), one examines the error decomposition
@@
$+$ $\alpha_{m}^{2}(\|P_{\rho}^{1/2}e^{(m)}\|^{2}-2\mu_{m}(P_{\rho}e^{(m)},P_{\omega_{m}}e^{(m)})+\mu_{m}^{2}\|P_{\rho}^{1/2}P_{\omega_{m}}e^{(m)}\|^{2}).$
@@
$2\alpha_{m}^{2}\mu_{m}\|P_{\rho}e^{(m)}\|^{2}$
on the right-hand side, which must compensate for positive contributions from terms such as
@@
Since in general $P_{\rho}$ does not commute with the operators $P_{\omega}$, it is not clear how to relate these quantities without additional assumptions.
@@ 5.3 A divergence result
$g:=\mathbb{E}(R_{\omega}y)=\mathbb{E}(R_{\omega}f_{\mu}(\omega))\not\in\mathrm{ran}(P_{\rho}),$
then the sequence $u^{(m)}$ obtained from the online algorithm (3)-(5) diverges in expectation to $\infty$ in $V$ for any choice of the parameters $1/2<t<1$ and $0<A\leq(2\Lambda)^{-1}$. This negative result is equivalent to proving
@@
and shows that (10) is essential in Theorem 2.1. As before, norm and scalar product in $V$ are denoted by $\|\cdot\|$ and $(\cdot,\cdot)$, respectively.
To establish (27), we expand $g$ and $u^{(0)}$ with respect to the CONS $\Psi$ and derive an inhomogeneous linear recursion for the coefficients of the expected error trajectory $U^{(m)}:=\mathbb{E}(u^{(m)})$. To do this, set
@@
$U^{(m+1)}=\alpha_{m}(U^{(m)}+\mu_{m}(g-P_{\rho}U^{(m)}))=\alpha_{m}(I-\mu_{m}P_{\rho})U^{(m)}+\alpha_{m}\mu_{m}g,\quad m\geq 0,$
@@
$m=1,2,\ldots$. By linearity of the expectation operator and the fact that $\Psi$ consists of the eigenvectors of $P_{\rho}$, we get
@@
$m=1,2,\ldots$, separately for each $i\in\mathbb{N}$. Under the restrictions on the parameters $\mu_{m}$ from (5), we have $1/2\leq 1-\mu_{l}\lambda_{i}\leq 1$. This shows that we have $0<\epsilon_{m}(\lambda_{i})\leq 1$ for the factor in front of $c_{i}$, which implies that the contribution of the initial guess $u^{(0)}$ can be neglected if $m\to\infty$.
@@
$\geq$ $\frac{C_{0}^{2}}{2}\left(\sum_{\lambda_{i}\geq C_{1}(m+1)^{t-1}}\lambda_{i}^{-2}g_{i}^{2}\right)-\frac{\|u^{(0)}\|^{2}}{(m+1)^{2}},$
@@
It remains to show (29). For this, set
$\Pi_{k}^{m-1}(\lambda):=\prod_{l=k}^{m-1}(1-\mu_{k}\lambda),\quad k=0,\ldots,m-1,\qquad\Pi_{m}^{m-1}(\lambda):=1.$
@@
$d_{m}(\lambda)\geq\frac{1}{2}\sum_{k\in I_{m}(\lambda)}k\mu_{k-1},\qquad I_{m}(\lambda):=\{k\leq m:\;C\lambda((m+1)^{1-t}-(k+1)^{1-t})\leq 1\}.$ (30)
@@
When we substitute this, $k\geq k_{0}+1\geq C_{2}(m+1)$, and $\mu_{k-1}=Ak^{-t}$ into (30), we get
@@ 5.4 A special case
$y_{m}=(u,\psi_{i_{m}})+\epsilon_{m},\qquad m=0,1,\ldots$
@@
where the coefficients $\alpha_{m}$ and $\mu_{m}$ are given by (5) with $\Lambda=1$. This is a special instance of (3) if we set $\Omega=\mathbb{N}$, $Y=\mathbb{R}$ and define $R_{i}:\,\mathbb{R}\to V$ and $R^{\ast}_{i}:\,V\to\mathbb{R}$ by $R_{i}y=y\psi_{i}$ and $R_{i}^{\ast}v=(v,\psi_{i})$, respectively. The associated RKHS $H$ can be identified with $\ell^{2}(\mathbb{N})$. To simplify things further, let $i_{m}$ be i.i.d. samples from $\mathbb{N}$ with respect to a discrete probability measure $\rho$ on $\mathbb{N}$, and let $\epsilon_{m}$ be i.i.d. random noise with zero mean and finite variance $\sigma^{2}<\infty$ that is independent of $i_{m}$. The corresponding operator $P_{\rho}$ is given by
@@
its eigenvalues $\lambda_{i}=\rho_{i}$ are given by $\rho$, and it is trace class (w.l.o.g., we assume $\rho_{1}\geq\rho_{2}\geq\ldots$). The spaces $V_{P_{\rho}}^{s}$, $\infty<s<\infty$, can now be identified as sets of formal orthogonal series
@@
and has $u$ as its unique solution. This example also shows that sometimes it is natural to consider convergence in $V$ rather than convergence in $L_{\rho}^{2}(\Omega,Y)$.
The simplicity of this example allows a comprehensive convergence theory with respect to the scale of $V_{P_{\rho}}^{s}$ spaces. We state the following results without proof.
@@
Let $1\leq\bar{s}\leq 0\leq s$, and $\bar{s}<s\leq\bar{s}+2$. Then, for the sampling process described above, the online algorithm (31) converges for $u\in V_{P_{\rho}}^{s}$ in the $V_{P_{\rho}}^{\bar{s}}$ norm with the bound
$\mathbb{E}(\|e^{(m)}\|^{2}_{V_{P_{\rho}}^{\bar{s}}})\leq C(m+1)^{-\min((s-\bar{s})/(s+2),2/(\bar{s}+4))}(A^{\bar{s}-s}\|u\|^{2}_{V_{P_{\rho}}^{s}}+A^{2+\bar{s}}\sigma^{2}),$ (32)
$m=1,2,\ldots$, if the parameters $t$ and $A$ in (5) satisfy
@@
Let $1\leq\bar{s}\leq 0\leq s$, $\bar{s}<s$ and $\sigma>0$. For the online algorithm (31) we have
$\sup_{\rho}\sup_{u:\,\|u\|_{V_{P_{\rho}}^{s}}=1}(m+1)^{\min((s-\bar{s})/(2+s),2/(\bar{s}+4))}\mathbb{E}(\|e^{(m)}\|^{2}_{V_{P_{\rho}}^{\bar{s}}})\geq c>0,$
@@
$\Pi_{k}^{m-1}:=(1-\frac{a}{m^{t}})\cdot\ldots\cdot(1-\frac{a}{(k+1)^{t}}),\quad 0\leq k\leq m-1,\quad\Pi_{m}^{m-1}:=1,$
@@
A matching lower bound for $\varepsilon_{m,i}$ can be obtained by using a slightly different value of $a$ in the definition of the products $\Pi_{k}^{m-1}$. The remainder of the argument for Theorem 5.1 first requires the substitution of tight upper bounds for $\Pi_{k}^{m-1}$ and $S_{k}$ in dependence on $t$ and $a$ into the bounds for $\varepsilon_{m,i}$. Next, after substituting the estimates for $\varepsilon_{m,i}$ into (33), the resulting series has to be estimated separately for the index sets $I_{1}:=\{i:\,A\rho_{i}\leq(m+1)^{t-1}\}$ and $I_{2}:=\mathbb{N}\backslash I_{1}$ followed by choosing the indicated optimal value of $t=t_{s,\bar{s}}$. This leads to the bound (32) in Theorem 5.1. For the proof of Theorem 5.2, lower bounds are needed for $\Pi_{k}^{m-1}$, $S_{k}$, and consequently $\varepsilon_{m}$, combined with choosing suitable discrete probability distributions $\rho$. Regarding lower bounds for $\Pi_{k}^{m-1}$, see the considerations in Subsection 5.3.